Problem No. 2

From following date compute (i) Real GDP (ii) GDP deflator (iii) inflation rate 


Year

Nominal GDP

Rs. billion

CPI


2018-19

54195

228.1


2020-21

61583

251.3



1
Expert's answer
2021-04-19T18:51:30-0400

(i) Real GDP in 2020-21 is:

Yr = 61583×228.1/251.3 = 55897.66.

(ii) GDP deflator is:

Deflator = 61583/55897.66×100 = 110.17.

(iii) Inflation rate is 110.17 - 100 = 10.17%.